HELIUM AND HYDROGEN INDUCED ROTATIONAL RELAXATION OF H_{2}CO OBSERVED AT TEM-PERATURES OF THE INTERSTELLAR MEDIUM

\underline{MARKUS MENGEL} and FRANK C. DE LUCIA, Department of Physics, The Ohio State University, 174 West 18th Ave., Columbus OH 43210, USA.

We have performed pressure broadening and time resolved double resonance studies on three rotational transitions (2_{12} ← 1_{11}, 2_{11} ← 1_{10}, and 3_{13} ← 2_{12}) of formaldehyde (H_{2}CO) in collision with helium and hydrogen at low temperatures (T ≤ 16 K). This was achieved by applying the collisional cooling method. The purpose of this study was to compare laboratory measurements with the theory used to explain the anomalous absorption of interstellar formaldehyde against the 2.7 K background.^{a} We can confirm this prediction as far as it constitutes a model based on H_{2}CO - He collisions. However, we observe significantly different pressure broadening cross sections for H_{2}CO - H_{2} collisions which raises the question as to whether a simple transfer from helium to hydrogen as collision partner can be made.

aB.J. Garrison, W.A. Lester Jr., and W.H. Miller, J. Chem. Phys. 65, 2193-2200 (1976)